OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2x - 2 <= 4? 3 >= x x <= 3 A number line with a closed circle on 3 and shading to the right. A number line with a closed circle on 3 and shading to the left.

OpenStudy (anonymous):

@TRowstudy

OpenStudy (anonymous):

@ovenmitt12

OpenStudy (richyw):

\[2x-2\leq 4\]\[2x\leq 6\]\[x\leq 3\]

OpenStudy (richyw):

\(a<x\) is the same thing as \(x>a\)

OpenStudy (anonymous):

The purpose of this is to show 1) that you understand that the first two are the same and 2) to show that you understand what that means on a number line. 3>=x is the same as x<=3. Now you know both of those must be correct, so one of the correct answers is the third or fourth options. You shade a number line with the values that satisfy the inequality. A number line with a closed circle on 3 and shading to the right would be the same thing as x>=3. A number line with a closed circle on 3 and shading to the left would mean x<=3.

OpenStudy (richyw):

you would draw a closed circle because " \(\leq\) " means "less than OR equal"

OpenStudy (richyw):

so you want to shade everything that it less than or equal to 3. I think you can figure out which one is incorrect from there!